# MOOSE-Chem3: Toward Experiment-Guided Hypothesis Ranking via Simulated Experimental Feedback

### v2

@@ Appendix B The Role of CriticalPoints in CSX-Sim
Final Correction Factor 0 End
The our simulator analyzes Hypothesis 4, noting that enhancing the Seebeck coefficient of a polymer gel material requires a redox pair. The first three hypotheses incorporate redox pairs, either [Fe(CN)6]3-/[Fe(CN)6]4- or Fe3+/Fe2+, which facilitate efficient conversion of thermal energy to electrical energy. In contrast, Hypothesis 4 only involves polymerizing acrylamide (PAM) and lacks a redox pair, rendering it unable to effectively convert thermal energy into electrical energy. Consequently, the thermoelectric potential (Seebeck coefficient, Se) is zero.
@@ C.1 Dataset Composition and Analysis
To evaluate the performance of the simulator, we conducted a thorough analysis using real-world experimental data. We curated a set of 30 cutting-edge chemical questions, each designed to probe significant aspects of chemical research. These questions were carefully selected to encompass multiple areas within the chemistry domain, ensuring a diverse and representative evaluation framework. Each question was associated with 3 to 6 hypotheses, resulting in a total of 124 authentic wet lab chemical experiment results. This extensive dataset forms a robust foundation for assessing the simulator’s predictive accuracy and reliability.
@@
Polymer Chemistry 16
Organic Chemistry 36
Inorganic Chemistry 33
Analytical Chemistry 39
@@
Table 5: Distribution of categories.
@@ C.2 Trend Comparison with Real Experiment Results
Figure 6 compares the trends observed in CSX-Sim predictions (green line) with those from real experimental data (blue line). Error bars, representing the population standard deviation, illustrate the variability of the data points. Statistical significance was further established using the Bootstrap method, with results indicating ($p<0.01$(Berg-Kirkpatrick et al., 2012). The aggregated analysis reveals that the simulator effectively predicts the mean trends for all 30 sets of results, demonstrating a strong consistency with the mean of the actual experimental outcomes. This alignment of mean trends across the diverse questions underscores the simulator’s ability to model chemical processes accurately, capturing the overall behavior of the experimental data, regardless of the specific subfield.
@@
Appendix D Evaluation of Trend Alignment and Accuracy
D.1 Evaluation of Trend Alignment
@@
Where: $d_{i}$: The difference between the ranks of the $i$-th simulated and experimental result. $n$: The number of hypotheses in a given group (ranging from 3 to 6 per scientific question). $\rho$: The correlation coefficient, ranging from -1 (perfect negative correlation) to 1 (perfect positive correlation), with 0 indicating no monotonic relationship. A Spearman Correlation Coefficient ($\rho$) near 1 indicates strong trend alignment, meaning the simulated results closely mirror the relative ordering of experimental outcomes. Our CSX simulator achieved a mean Spearman Correlation Coefficient of $\rho=0.960$, significantly outperforming the baseline, as shown in Table 1, and demonstrating superior trend alignment.
To further assess the robustness of the simulator across diverse problems, we introduced the Perfect Consistency Indicator (PCI), a stringent metric that counts the number of question groups (out of the 30 scientific questions) where the simulated results achieved perfect trend alignment with the experimental results $(\rho=1)$. Perfect trend alignment requires an exact match in the ranking of simulated and experimental outcomes, making PCI a robust measure of the simulator’s ability to consistently replicate experimental trends across all problems. Notably, our CSX simulator achieved perfect trend alignment $(\rho=1)$ in 26 out of 30 question groups, significantly surpassing the baseline methods and highlighting its exceptional robustness and predictive fidelity.
D.2 Evaluation of Simulator Accuracy
@@
Where: $y_{i}$: The experimental result for the $i$-th hypothesis. $\hat{y}_{i}$: The simulated result for the $i$-th hypothesis. The CSX simulator exhibited a lower RMSE than the "Matched Score" baseline (Yang et al., 2024b), signifying improved predictive accuracy, as substantiated by the results in Table 1.
To thoroughly evaluate the predictive accuracy of our CSX simulator compared to real-world experimental outcomes, we tested its performance on a dataset of 124 authentic scientific hypotheses. For a comprehensive comparison, we calculated several performance indicators, as presented in Table 6. Building on the previously discussed metrics, we introduced three additional measures: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Logarithmic Error (RMSLE). These metrics, defined below, enhance the robustness of our analysis by capturing different aspects of prediction error.
@@
As shown in Table 6, CSX-Sim consistently outperforms the "Matched Score" baseline (Yang et al., 2024b) across all metrics, achieving an MSE of 0.058, an MAE of 0.161, and an RMSLE of 0.147. Ablation studies further reveal the contributions of individual components: the removal of CriticalPoints results in a slight performance decline (MSE of 0.064, MAE of 0.174, RMSLE of 0.159), while the exclusion of ComponentExtraction leads to more significant degradation (MSE of 0.087, MAE of 0.215, RMSLE of 0.192). These results underscore the importance of both critical point identification and component extraction in achieving high predictive accuracy and robustness in simulation outcomes.
Appendix E Different Levels of Distortion
@@
Appendix F Evaluation of Experiment-Guided Ranking and Its Societal Benefits
@@
Appendix G Limitations
A primary limitation of this work is that the constructed simulator does not provide perfectly accurate experimental feedback. Specifically, the simulator is based on three foundational assumptions developed through extensive consultations with domain experts. While it represents the first attempt to build such a simulator for experiment-guided hypothesis ranking, its outputs remain an approximation rather than exact experimental results.
The rationale for developing this simulator stems from the absence of any prior tools with comparable functionality. Its purpose is to enable research on experiment-guided ranking methods, which can later be applied and validated with real experimental feedback in practical settings.
Importantly, the simulator’s absolute accuracy is not critical for this line of research. As long as the experiment-guided ranking methods are robustly developed and tested within this simulated environment, they can subsequently leverage real experimental feedback to identify optimal hypotheses when deployed in real-world scenarios.

### v3

@@ Appendix B The Role of CriticalPoints in CSX-Sim
Final Correction Factor 0 Endo
Our simulator analyzes Hypothesis 4, noting that enhancing the Seebeck coefficient of a polymer gel material requires a redox pair. The first three hypotheses incorporate redox pairs, either [Fe(CN)6]3-/[Fe(CN)6]4- or Fe3+/Fe2+, which facilitate efficient conversion of thermal energy to electrical energy. In contrast, Hypothesis 4 only involves polymerizing acrylamide (PAM) and lacks a redox pair, rendering it unable to effectively convert thermal energy into electrical energy. Consequently, the thermoelectric potential (Seebeck coefficient, Se) is zero.
@@ C.1 Dataset Composition and Analysis
To evaluate the performance of the simulator, we conducted a thorough analysis using real-world experimental data. We curated a set of 30 cutting-edge research questions, each designed to probe significant aspects of chemical research. These questions were carefully selected to encompass multiple areas within the chemistry domain, ensuring a diverse and representative evaluation framework. Each question was associated with 3 to 6 hypotheses, resulting in a total of 124 authentic wet lab experiment results. This extensive dataset forms a robust foundation for assessing the simulator’s predictive accuracy and reliability.
@@
Energy Materials 12
Polymeric Materials 8
Applied Physics 18
Systems Biology 10
Organic Chemistry 26
Inorganic Chemistry 24
Analytical Chemistry 26
@@
Table 5: Classification of the 124 real-world experiments used to validate the simulator.
@@ C.2 Trend Comparison with Real Experiment Results
Figure 6 compares the trends observed in CSX-Sim predictions (green line) with those from real experimental data (blue line). Error bars, representing the population standard deviation, illustrate the variability of the data points. Statistical significance was further established using the Bootstrap method, with results indicating (p<0.01p<0.01(Berg-Kirkpatrick et al., 2012). The aggregated analysis reveals that the simulator effectively predicts the mean trends for all 30 sets of results, demonstrating a strong consistency with the mean of the actual experimental outcomes. This alignment of mean trends across the diverse questions underscores the simulator’s ability to model chemical processes accurately, capturing the overall behavior of the experimental data, regardless of the specific subfield.
@@
Appendix D Disciplines of the TOMATO-chem benchmark dataset
Category Count
Energy Materials 363
Polymeric Materials 359
Metallic Materials 268
Nanomaterials 316
Applied Physics 127
Analytical Chemistry 317
Inorganic Chemistry 392
Organic Chemistry 482
Chemical Engineering 196
Environmental Engineering 298
Molecular Biology 84
Systems Biology 62
Total 3264
Table 6: Disciplinary classification of the TOMATO-chem benchmark dataset
Appendix E Evaluation of Trend Alignment and Accuracy
E.1 Evaluation of Trend Alignment
@@
Where: $d_{i}$: The difference between the ranks of the $i$-th simulated and experimental result. $n$: The number of hypotheses in a given group (ranging from 3 to 6 per scientific question). $\rho$: The correlation coefficient, ranging from -1 (perfect negative correlation) to 1 (perfect positive correlation), with 0 indicating no monotonic relationship. A Spearman Correlation Coefficient ($\rho$) near 1 indicates strong trend alignment, meaning the simulated results closely mirror the relative ordering of experimental outcomes. Our CSX-Sim achieved a mean Spearman Correlation Coefficient of $\rho=0.960$, significantly outperforming the baseline, as shown in Table 1, and demonstrating superior trend alignment.
To further assess the robustness of the simulator across diverse problems, we introduced the Perfect Consistency Indicator (PCI), a stringent metric that counts the number of question groups (out of the 30 scientific questions) where the simulated results achieved perfect trend alignment with the experimental results $(\rho=1)$. Perfect trend alignment requires an exact match in the ranking of simulated and experimental outcomes, making PCI a robust measure of the simulator’s ability to consistently replicate experimental trends across all problems. Notably, our CSX-Sim achieved perfect trend alignment $(\rho=1)$ in 26 out of 30 question groups, significantly surpassing the baseline methods and highlighting its exceptional robustness and predictive fidelity.
E.2 Evaluation of Simulator Accuracy
@@
Where: $y_{i}$: The experimental result for the $i$-th hypothesis. $\hat{y}_{i}$: The simulated result for the $i$-th hypothesis. The CSX-Sim exhibited a lower RMSE than the "Matched Score" baseline (Yang et al., 2025), signifying improved predictive accuracy, as substantiated by the results in Table 1.
To thoroughly evaluate the predictive accuracy of our simulator compared to real-world experimental outcomes, we tested its performance on a dataset of 124 authentic scientific hypotheses. For a comprehensive comparison, we calculated several performance indicators, as presented in Table 7. Building on the previously discussed metrics, we introduced three additional measures: Mean Squared Error (MSE), Mean Absolute Error (MAE), and Root Mean Squared Logarithmic Error (RMSLE). These metrics, defined below, enhance the robustness of our analysis by capturing different aspects of prediction error.
@@
As shown in Table 7, CSX-Sim consistently outperforms the "Matched Score" baseline (Yang et al., 2025) across all metrics, achieving an MSE of 0.058, an MAE of 0.161, and an RMSLE of 0.147. Ablation studies further reveal the contributions of individual components: the removal of CriticalPoints results in a slight performance decline (MSE of 0.064, MAE of 0.174, RMSLE of 0.159), while the exclusion of ComponentExtraction leads to more significant degradation (MSE of 0.087, MAE of 0.215, RMSLE of 0.192). These results underscore the importance of both critical point identification and component extraction in achieving high predictive accuracy and robustness in simulation outcomes.
Appendix F Different Levels of Distortion
@@
Appendix G Performance comparison of different functions in the simulator
Table 9: Performance comparison of different functions in the simulator.
Function Spearman Corr. ($\uparrow$) RMSE ($\downarrow$) Perfect Consistency ($\uparrow$)
Linear Function 0.9708 0.1959 24/30
Gaussian Function 0.9600 0.2147 26/30
Absolute Value Function 0.9626 0.2595 23/30
Quadratic Function 0.9682 0.3996 22/30
This supplementary study was conducted to validate the robustness of our core mathematical modeling. While the Gaussian function was selected for its well-behaved mathematical properties and intuitive alignment with our core assumptions, the framework’s success is not tied to any single function form. The choice of function can be viewed as a tunable hyperparameter.
Table 9 presents the results of a comparative study of various monotonic functions serving as the core of the simulator. Performance was assessed on Spearman Correlation ($\uparrow$), RMSE ($\downarrow$), and Perfect Consistency ($\uparrow$). The results show that all tested functions provide effective ranking guidance, which underscores the framework’s overall robustness. This analysis confirms that our framework is adaptable and can accommodate different function forms, enhancing its generalizability across domains.
Appendix H Evaluation of Experiment-Guided Ranking and Its Societal Benefits
@@
Appendix I Scalability Analysis
Method Trials (N = 64) Trials (N = 128)
Uninformed Search 32.5 64.5
Pre-experiment ranking 28.6 51.3
CSX-Rank (ours) 15.2 30.7
To evaluate the scalability of our proposed method, we expanded the pool of candidate hypotheses from N=64 to N=128. The results, presented in Table  10, show that our method, CSX-Sim, required 15.2 trials for 64 candidates and 30.7 trials for 128 candidates. In contrast, Uninformed Search and Pre-experiment ranking required 64.5 and 51.3 trials, respectively, for the larger candidate pool. The experimental results align with our theoretical analysis. The performance of CSX-Rank demonstrates a near-linear growth, which is consistent with its average-case time complexity of $O(N)$. The observed slope of approximately $0.24\times N$ confirms this scalability and shows that CSX-Rank retains a substantial cost advantage even as the candidate pool expands. Theoretically, in a best-case scenario where the clustering of hypotheses is highly effective, the cost could be reduced to $O(\log N)$, as evidence from one experiment can be generalized to every hypothesis within its cluster.